# On the kernel of $\mathrm{SO}(3)$-Witten-Reshetikhin-Turaev quantum representations

### v2

@@ 1. Introduction
Our last result concerns the kernel of $\rho_{p,2}$. In [GM14], a decomposition of the representations $\rho_{p,1}$ as the sum of two irreducible representations of $\operatorname{\mathrm{Mod}}(\Sigma)$ over $\mathbb{F}_{p}$ is introduced (which is called the odd-even decomposition by Gilmer and Masbaum). Using this decomposition and the structure of the abelianization of the Johnson subgroup $J_{2}(\Sigma)$ for surfaces $\Sigma$ of genus $g\geq 6,$ and the odd-even decomposition of $\rho_{p},$ we prove:
@@ 2.1. Basic properties of $\mathrm{SO}(3)$-WRT representations
In this section, we will sketch a construction of the representation $\rho_{p},$ and state some of its properties. Let $p$ be an odd integer $\geq 3,$ let $M$ be a closed compact oriented $3$-manifold, and let $L$ be a framed link in $M.$ In [BHMV92], a topological invariant $Z_{p}(M,L)\in\mathbb{Q}[\zeta_{p}]$ of the pair $(M,L)$ is defined. Roughly speaking, $Z_{p}(M,L)$ is the evaluation at a primitive $2$p-th root of unity $\zeta_{p}$ of a suitable linear combination of the colored Jones polynomials of the link $L_{0}\cup L,$ where $L_{0}$ is a surgery presentation of $M.$ Here suitable means that the invariant thus defined is independent of the choice of surgery presentation $L_{0}$ for $M,$ this is achieved by coloring the surgery presentation $L_{0}$ by the so-called Kirby color. We refer to [BHMV92] for all details on this construction; we will use only some of the properties of this invariant $Z_{p}(M,L).$ First we note that as a consequence of the construction, the invariant $Z_{p}(M,L)$ satisfies the Kauffman relations in terms of the link $L.$ Moreover, by a theorem of Masbaum and Roberts [MR97], when $p$ is prime then $Z_{p}(M,L)\in\mathbb{Z}[\zeta_{p}].$
@@
Given a lollipop tree decomposition of $\Sigma,$ we associate elements $v(a,b)\in\mathcal{S}_{p}(\Sigma)$ to colorings of the edges of the trivalent graph $\Gamma$ associated to the decomposition as follows. We take $H$ to be a handlebody with $\partial H=\Sigma,$ and such that the curve of the pants decomposition bound disks in $H.$ We let the internal edges of $\Gamma$ be colored by $(2a_{i})_{1\leq i\leq 2g-3},$ and we assume that $2a_{1},\ldots,2a_{g}$ are the colors of the edges that are adjacent to one-edge loops. The loop edges are colored by $a_{i}+b_{i}.$ We assume again that if $c_{1},c_{2},c_{3}$ are the colors around a trivalent vertex, then we have the $p$-admissibility conditions: $c_{1}\leq c_{2}+c_{3},$ and $c_{1}+c_{2}+c_{3}$ is even and $\leq 2p-4.$ Moreover, we ask for the colors $a_{i}+b_{i}$ to be at most $(p-3)/2,$ which Gilmer and Masbaum call a small coloring. We construct an element $\Gamma(a,b)\in\mathcal{S}_{p}(\Sigma)$ by embedding the trivalent graph $\Gamma$ in $H,$ replacing the internal edges with $2a_{i}$ parallel copies with the Jones-Wenzl idempotent $f_{2a_{i}}$ inserted. However, for the loop edge colored by $a_{i}+b_{i},$ we first cable the loop edge by the Jones-Wenzl idempotent $f_{a_{i}},$ then we add a copy of the loop colored by $(\frac{2+z}{h})^{b_{i}},$ where $h=1-\zeta_{p}.$ (Here we use the convention that a framed link colored by $z^{n}$ stands for $n$ parallel copies of the link, and we can make sense of framed links colored by a poloynomial in $z$ by extending linearly). Finally, we define $v(a,b)=h^{-\lfloor\frac{1}{2}(a_{1}+\ldots+a_{g})\rfloor}\Gamma(a,b).$
@@
for some $A\in\mathrm{End}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma)).$
@@
for some $A_{1}\in\mathrm{End}(\mathcal{S}_{p}^{odd}(\Sigma)),$ $A_{2}\in\mathrm{End}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma))$ and $A_{3}\in\mathrm{End}(\mathcal{S}_{p}^{ev}(\Sigma)).$
@@
for some $B\in\mathrm{End}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma)).$
@@ 2.2. Computations of $\rho_{p}$ on bounding pairs and separating twists
Next we prove a linear independence result for the $\rho_{p,2}$ image of some separating twists. Recall that $\mathrm{Im}\rho_{p,2}|_{J_{2}(\Sigma)}$ has a natural structure of $\mathbb{F}_{p}$-vector space, obtained by considering the order $1$ in $h.$ (see also Lemma 3.4)
@@
In the basis $v(a,b),$ we will denote by $a_{i}$ the color corresponding to the curve $\alpha_{i}.$ For $1\leq i\leq 2g-3,$ let $v_{i}$ be the vector $v(a,0),$ where $a_{i}=2$ and $a_{j}=1$ for $j\neq i.$ Let also $w$ be the vector $v(a,0),$ where $a_{i}=1$ for all $i.$ The separating Dehn twists $t_{\alpha_{i}}$ acts diagonally on the basis $v(a,b),$ and moreover we have:
$\rho_{p}(t_{\alpha_{i}})(v(a,b))=\zeta_{p}^{2a_{i}(2a_{i}+2)}v(a,b),$
@@
$\rho_{p,2}(t_{\alpha_{i}})(v_{j})=\begin{cases}(1+24h)v_{i}\ (\mathrm{mod}\ h^{2})\ \textrm{if }i=j\\ (1+8h)v_{j}\ (\mathrm{mod}\ h^{2})\ \textrm{else}\end{cases}$
@@
$8\begin{pmatrix}3&1&\ldots&1\\ 1&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&1\\ 1&\ldots&1&3\end{pmatrix}$
has determinant $8^{2g-3}2^{2g-4}(2g-1),$ which is non-zero mod $p$ unless $p$ divides $12g-$ 1 . In the latter case its kernel is generated by (${}^{t}(1,1,\ldots,1).$ However, in that case we also compute $\rho_{p,2}(t_{\alpha_{1}}\ldots t_{\alpha_{2g-3}})(w)=(1+8(2g-3)h)w\ (\mathrm{mod}\ h^{2}),$ hence the $\rho_{p,2}(t_{\alpha_{i}})$ are still linearly independent. ∎
Recall that if $\Sigma$ is a surface of genus at least two with one boundary component, we denote by $\hat{\Sigma}$ the surface obtained by gluying a disk on the boundary of $\Sigma$. We also denote by $\rho_{p}^{s}$ the representation of $\pi_{1}(\hat{\Sigma})$ obtained by restriction of $\rho_{p}$ where the boundary color is $2$.
@@ 2.4. Structure of the abelianization of the Torelli and Johnson subgroups
The result of Johnson actually explicitely describes the $\mathrm{Sp}_{2g}(\mathbb{Z})$-module structure of the $2$-torsion group, but we will not need it here; we will only use that it is a $2$-torsion group.
@@
    The factor $\mathbb{Q}$ in the above isomorphism is the rational span of the core of the Casson invariant, which by [Mor91, Theorem 5.7] is a morphism
@@
    The factor $[2^{2}]$ is the rational image of the second Johnson homomorphism, which is the map
@@
    The factor $[31^{2}]$ is isomorphic to $(J_{3}/J_{4})_{\mathbb{Q}}:=\left(J_{3}(\Sigma)/J_{4}(\Sigma)\right)\otimes\mathbb{Q}$ as a $\mathrm{Sp}_{2g}(\mathbb{Z})$-representation. It is non-canonical, and depends on a choice of splitting of the exact sequence
@@
    and thus an isomorphism $(J_{2}/J_{4})_{\mathbb{Q}}\simeq(J_{2}/J_{3})_{\mathbb{Q}}\times(J_{3}/J_{4})_{\mathbb{Q}}.$ (we remark that $J_{2}(\Sigma)/J_{4}(\Sigma)$ is abelian since $[J_{2}(\Sigma),J_{2}(\Sigma)]\subset J_{4}(\Sigma).$) The map $\mathrm{Ab}(J_{2}(\Sigma))\longrightarrow[2^{2}]\oplus[31^{2}]$ is the so-called refined second Johnson homomorphism, which was first introduced in [Mor99] and [MSS15]. This morphism vanishes on $J_{4}(\Sigma),$ as noted in the second equation of [MSS20] between Equations (9) and (10), and as follows from our description. Finally, it can be noted that the map $\mathrm{Ab}(J_{2}(\Sigma))\longrightarrow[31^{2}]$ does not vanish on $[J_{1}(\Sigma),J_{2}(\Sigma)]$ (as we shall also see from our computations).
@@ 2.5. Irreducibility of some modular representations of $\mathrm{Sp}_{2g}(\mathbb{Z})$
On the other direction, let $W$ be a subrepresentation of $V.$ Let $c_{1},\ldots,c_{2g}=a_{1},b_{1},\ldots,a_{g},b_{g}$ be a symplectic basis of $H_{1}(\Sigma,\mathbb{F}_{p}).$ We also write $c_{i}^{\prime}=c_{i-1}$ if $i$ is even, and $c_{i}^{\prime}=c_{i+1}$ if $i$ is odd, so that $\{c_{i},c_{i}^{\prime}\}$ is always a pair $\{a_{k},b_{k}\}.$ We note that the vectors $c_{i}\wedge c_{j}\wedge c_{k}$ for $i<j<k$ and $(i,j,k)\neq(i,2g-1,2g)$ and $(2g-3,2g-2,k)$ form a basis of $V.$
@@ 3.3. Analysis of $\mathrm{Ker}(\rho_{p,2})$
The inclusion $[J_{2}(\Sigma),J_{2}(\Sigma)]T_{p}\subset\ker\rho_{p,2}$ is clear, since $T_{p}\subset\ker\rho_{p},$ and $\rho_{p,2}$ restricted to $J_{2}(\Sigma)$ has abelian image by Lemma 3.4. By Theorem 1.2, we have that $\operatorname{\mathrm{Ker}}\rho_{p,1}\subset[J_{1}(\Sigma),J_{1}(\Sigma)]T_{p}\subset J_{2}(\Sigma)T_{p}.$ Since $T_{p}\subset\operatorname{\mathrm{Ker}}\rho_{p},$ we only need to describe which elements of $J_{2}(\Sigma)$ are in $\operatorname{\mathrm{Ker}}\rho_{p,2}.$ However, by Proposition 3.5, for $p$ large enough, we have that $f\in J_{2}(\Sigma)$ is in $\operatorname{\mathrm{Ker}}\rho_{p,2}$ if and only if $f$ is in the kernel of the mod $p$ abelianization morphism. Since $J_{2}(\Sigma)$ is generated by separating twists, the kernel of the mod $p$ abelianization is generated by commutators in $J_{2}(\Sigma)$ and $p$-th powers of separating Dehn twists, both of which are in $[J_{2}(\Sigma),J_{2}(\Sigma)]T_{p}.$
@@ 4. Further comments
Lemma 4.1 makes it tempting to think of the morphism $d$ as (a scalar multiple of) the reduction mod $p$ of the core of the Casson invariant, which is the unique (up to scalar) $\mathrm{Mod}(\Sigma)$-invariant morphism $J_{2}(\Sigma)\rightarrow\mathbb{F}_{p}$ by the work of Morita [Mor89]. Unfortunately, nothing garanties that the morphism $d$ is not identically zero.
To check whether this is the case, it is possible to compute $d$ on a separating Dehn twist of genus $1,$ thanks to Gilmer and Masbaum’s formulae for the dimensions of the odd-even decomposition of $\mathcal{S}_{p}(\Sigma)$ (see [GM14, Proposition 7.2]). By performing this analysis, it is possible to check that for each genus $g,$ the morphism $d^{\prime}$ is non trivial for small values of $p,$ and trivial for all large enough primes $p.$ Since the proof of this negative result is a bit involved, we will not include it here.

### v3

@@ 1. Introduction
Our last result concerns the kernel of $\rho_{p,2}$. In [GM14], a decomposition of the representations $\rho_{p,1}$ as the sum of two irreducible representations of $\operatorname{\mathrm{Mod}}(\Sigma)$ over $\mathbb{F}_{p}$ is introduced (which is called the odd-even decomposition by Gilmer and Masbaum). Using the structure of the abelianization of the Johnson subgroup $J_{2}(\Sigma)$ for surfaces $\Sigma$ of genus $g\geq 6,$ and the odd-even decomposition of $\rho_{p},$ we prove:
@@ 2.1. Basic properties of $\mathrm{SO}(3)$-WRT representations
In this section, we will sketch a construction of the representation $\rho_{p},$ and state some of its properties. Let $p$ be an odd integer $\geq 3,$ let $M$ be a closed compact oriented $3$-manifold, and let $L$ be a framed link in $M.$ In [BHMV92], a topological invariant $Z_{p}(M,L)\in\mathbb{Q}[\zeta_{p}]$ of the pair $(M,L)$ is defined. Roughly speaking, $Z_{p}(M,L)$ is the evaluation at a primitive $p$-th root of unity $\zeta_{p}$ of a suitable linear combination of the colored Jones polynomials of the link $L_{0}\cup L,$ where $L_{0}$ is a surgery presentation of $M.$ Here suitable means that the invariant thus defined is independent of the choice of surgery presentation $L_{0}$ for $M,$ this is achieved by coloring the surgery presentation $L_{0}$ by the so-called Kirby color. We refer to [BHMV92] for all details on this construction; we will use only some of the properties of this invariant $Z_{p}(M,L).$ First we note that as a consequence of the construction, the invariant $Z_{p}(M,L)$ satisfies the Kauffman relations in terms of the link $L.$ Moreover, by a theorem of Masbaum and Roberts [MR97], when $p$ is prime then $Z_{p}(M,L)\in\mathbb{Z}[\zeta_{p}].$
@@
Given a lollipop tree decomposition of $\Sigma,$ we associate elements $v(a,b)\in\mathcal{S}_{p}(\Sigma)$ to colorings of the edges of the trivalent graph $\Gamma$ associated to the decomposition as follows. We take $H$ to be a handlebody with $\partial H=\Sigma,$ and such that the curve of the pants decomposition bound disks in $H.$ We let the internal edges of $\Gamma$ be colored by $(2a_{i})_{1\leq i\leq 2g-3},$ and we assume that $2a_{1},\ldots,2a_{g}$ are the colors of the edges that are adjacent to one-edge loops. The loop edges are colored by $a_{i}+b_{i}.$ We assume again that if $c_{1},c_{2},c_{3}$ are the colors around a trivalent vertex, then we have the $p$-admissibility conditions: $c_{1}\leq c_{2}+c_{3},$ and $c_{1}+c_{2}+c_{3}$ is even and $\leq 2p-4.$ Moreover, we ask for the colors $a_{i}+b_{i}$ to be at most $(p-3)/2,$ which Gilmer and Masbaum call a small coloring. We construct an element $\Gamma(a,b)\in\mathcal{S}_{p}(\Sigma)$ by embedding the trivalent graph $\Gamma$ in $H,$ replacing the internal edges with $2a_{i}$ parallel copies with the Jones-Wenzl idempotent $f_{2a_{i}}$ inserted. However, for the loop edge colored by $a_{i}+b_{i},$ we first cable the loop edge by the Jones-Wenzl idempotent $f_{a_{i}},$ then we add a copy of the loop colored by $(\frac{2+z}{h})^{b_{i}},$ where $h=1-\zeta_{p}.$ (Here we use the convention that a framed link colored by $z^{n}$ stands for $n$ parallel copies of the link, and we can make sense of framed links colored by a polynomial in $z$ by extending linearly). Finally, we define $v(a,b)=h^{-\lfloor\frac{1}{2}(a_{1}+\ldots+a_{g})\rfloor}\Gamma(a,b).$
@@
for some $A\in\mathrm{Hom}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma)).$
@@
for some $A_{1}\in\mathrm{End}(\mathcal{S}_{p}^{odd}(\Sigma)),$ $A_{2}\in\mathrm{Hom}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma))$ and $A_{3}\in\mathrm{End}(\mathcal{S}_{p}^{ev}(\Sigma)).$
@@
for some $B\in\mathrm{Hom}(\mathcal{S}_{p}^{ev}(\Sigma),\mathcal{S}_{p}^{odd}(\Sigma)).$
@@ 2.2. Computations of $\rho_{p}$ on bounding pairs and separating twists
Next we prove a linear independence result for the $\rho_{p,2}$ image of some separating twists. Recall that $\mathrm{Im}\rho_{p,2}|_{J_{2}(\Sigma)}$ has a natural structure of $\mathbb{F}_{p}$-vector space, obtained by considering the order $1$ in $h$ (see also Lemma 3.4).
@@
In the basis $v(a,b),$ we will denote by $2a_{i}$ the color corresponding to the curve $\alpha_{i}.$ For $1\leq i\leq 2g-3,$ let $v_{i}$ be the vector $v(a,0),$ where $a_{i}=2$ and $a_{j}=1$ for $j\neq i.$ Let also $w$ be the vector $v(a,0),$ where $a_{i}=1$ for all $i.$
The separating Dehn twists $t_{\alpha_{i}}$ acts diagonally on the basis $v(a,b),$ and moreover we have:
$\rho_{p}(t_{\alpha_{i}})(v(a,b))=\zeta_{p}^{2a_{i}(a_{i}+1)}v(a,b),$
@@
$\rho_{p,2}(t_{\alpha_{i}})(v_{j})=\begin{cases}(1+12h)v_{i}\ (\mathrm{mod}\ h^{2})\ \textrm{if }i=j\\ (1+4h)v_{j}\ (\mathrm{mod}\ h^{2})\ \textrm{else}\end{cases}$
@@
$4\begin{pmatrix}3&1&\ldots&1\\ 1&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&1\\ 1&\ldots&1&3\end{pmatrix}$
has determinant $4^{2g-3}2^{2g-4}(2g-1),$ which is non-zero mod $p$ unless $p$ divides $12g-$ 1 . In the latter case its kernel is generated by (${}^{t}(1,1,\ldots,1).$ However, in that case we also compute $\rho_{p,2}(t_{\alpha_{1}}\ldots t_{\alpha_{2g-3}})(w)=(1+4(2g-3)h)w\ (\mathrm{mod}\ h^{2}),$ hence the $\rho_{p,2}(t_{\alpha_{i}})$ are still linearly independent.
Recall that if $\Sigma$ is a surface of genus at least two with one boundary component, we denote by $\hat{\Sigma}$ the surface obtained by gluing a disk on the boundary of $\Sigma$. We also denote by $\rho_{p}^{s}$ the representation of $\pi_{1}(\hat{\Sigma})$ obtained by restriction of $\rho_{p}$ where the boundary color is $2$.
@@ 2.4. Structure of the abelianization of the Torelli and Johnson subgroups
We note that the map
$\tau_{1}:\mathrm{Ab}(J_{1}(\Sigma))\longrightarrow\Lambda^{3}H_{1}(\Sigma,\mathbb{Z})/\left(\omega\wedge H_{1}(\Sigma,\mathbb{Z})\right)$
induced by the above isomorphism is also known as the first Johnson isomrphism, and the right hand-side is also isomorphic to $J_{1}(\Sigma)/J_{2}(\Sigma).$
Moreover, the result of Johnson actually explicitely describes the $\mathrm{Sp}_{2g}(\mathbb{Z})$-module structure of the $2$-torsion group, but we will not need it here; we will only use that it is a $2$-torsion group.
@@
    The summand $\mathbb{Q}$ in the above isomorphism is the rational span of the core of the Casson invariant, which by [Mor91, Theorem 5.7] is a morphism
@@
    The summand $[2^{2}]$ is the rational image of the second Johnson homomorphism, which is the map
@@
    The summand $[31^{2}]$ is isomorphic to $(J_{3}/J_{4})_{\mathbb{Q}}:=\left(J_{3}(\Sigma)/J_{4}(\Sigma)\right)\otimes\mathbb{Q}$ as a $\mathrm{Sp}_{2g}(\mathbb{Z})$-representation. The projection from $\mathrm{Ab}(J_{2}(\Sigma))$ to this summand is non-canonical, and depends on a choice of splitting of the exact sequence
@@
    and thus an isomorphism $(J_{2}/J_{4})_{\mathbb{Q}}\simeq(J_{2}/J_{3})_{\mathbb{Q}}\times(J_{3}/J_{4})_{\mathbb{Q}}.$ We remark that $J_{2}(\Sigma)/J_{4}(\Sigma)$ is abelian since $[J_{2}(\Sigma),J_{2}(\Sigma)]\subset J_{4}(\Sigma).$
    The map $\mathrm{Ab}(J_{2}(\Sigma))\longrightarrow[2^{2}]\oplus[31^{2}]$ is the so-called refined second Johnson homomorphism, which was first introduced in [Mor99] and [MSS15]. This morphism vanishes on $J_{4}(\Sigma),$ as noted in the second equation of [MSS20] between Equations (9) and (10), and as follows from our description. Finally, it can be noted that the map $\mathrm{Ab}(J_{2}(\Sigma))\longrightarrow[31^{2}]$ does not vanish on $[J_{1}(\Sigma),J_{2}(\Sigma)]$ (as we shall also see from our computations).
@@ 2.5. Irreducibility of some modular representations of $\mathrm{Sp}_{2g}(\mathbb{Z})$
Let $c_{1},\ldots,c_{2g}=a_{1},b_{1},\ldots,a_{g},b_{g}$ be a symplectic basis of $H_{1}(\Sigma,\mathbb{F}_{p}).$ We also write $c_{i}^{\prime}=c_{i-1}$ if $i$ is even, and $c_{i}^{\prime}=c_{i+1}$ if $i$ is odd, so that $\{c_{i},c_{i}^{\prime}\}$ is always a pair $\{a_{k},b_{k}\}.$ We note that the vectors $c_{i}\wedge c_{j}\wedge c_{k}$ for $i<j<k$ and $(i,j,k)\neq(i,2g-1,2g)$ and $(2g-3,2g-2,k)$ form a basis of $V.$
@@ 3.3. Analysis of $\mathrm{Ker}(\rho_{p,2})$
The inclusion $[J_{2}(\Sigma),J_{2}(\Sigma)]T_{p}\subset\ker\rho_{p,2}$ is clear, since $T_{p}\subset\ker\rho_{p},$ and $\rho_{p,2}$ restricted to $J_{2}(\Sigma)$ has abelian image by Lemma 3.4.
By Theorem 1.2, we have that $\operatorname{\mathrm{Ker}}\rho_{p,1}\subset[J_{1}(\Sigma),J_{1}(\Sigma)]T_{p}\subset J_{2}(\Sigma)T_{p}.$ Since $T_{p}\subset\operatorname{\mathrm{Ker}}\rho_{p},$ we only need to describe which elements of $J_{2}(\Sigma)$ are in $\operatorname{\mathrm{Ker}}\rho_{p,2}.$ However, by Proposition 3.5, for $p$ large enough, we have that $f\in J_{2}(\Sigma)$ is in $\operatorname{\mathrm{Ker}}\rho_{p,2}$ if and only if $f$ is in the kernel of the mod $p$ abelianization morphism. Since $J_{2}(\Sigma)$ is generated by separating twists, the kernel of the mod $p$ abelianization is generated by commutators in $J_{2}(\Sigma)$ and $p$-th powers of separating Dehn twists, both of which are in $[J_{2}(\Sigma),J_{2}(\Sigma)]T_{p}.$
@@ 4. Further comments
Lemma 4.1 makes it tempting to think of the morphism $d^{\prime}$ as (a scalar multiple of) the reduction mod $p$ of the core of the Casson invariant, which is the unique (up to scalar) $\mathrm{Mod}(\Sigma)$-invariant morphism $J_{2}(\Sigma)\rightarrow\mathbb{F}_{p}$ by the work of Morita [Mor89]. Unfortunately, nothing garanties that the morphism $d^{\prime}$ is not identically zero.
To check whether this is the case, it is possible to compute $d^{\prime}$ on a separating Dehn twist of genus $1,$ thanks to Gilmer and Masbaum’s formulae for the dimensions of the odd-even decomposition of $\mathcal{S}_{p}(\Sigma)$ (see [GM14, Proposition 7.2]). By performing this analysis, it is possible to check that for each genus $g,$ the morphism $d^{\prime}$ is non trivial for small values of $p,$ and trivial for all large enough primes $p.$ Since the proof of this negative result is a bit involved, we will not include it here.